Consider the following batch flow process consisting of four process steps operated by Anna, Aaron, Emily and Mike. Emily’s step requires set-up. Assume the units in a production batch must be moved together between steps. The activity times and setup time are given below: What batch size would you choose so as to maximize the process capacity?

a. 5

b. 10

c. 15

d. 20

e. 25

Process Steps operated by Processing Time in min per unit Setup time
Anna 6 0
Aaron 5 0
Emily 4 60
Mike 10 0
Answer #1

Let the required batch size = N

Total time taken by Emily to process batch size of N

= Processing time in minutes / unit x N units + Set up time

= 4.N + 60 minutes

Average time taken by 1 unit =( 4N + 60)/N = 4 + 60/N

Strictly based on processing time alone, the bottleneck activity is the one operated by Mike which has a processing time of 10 minutes per unit

Therefore in order to maximize process capacity , it has to be ensured that average time taken by Emily to process 1 unit shall not exceed 10 minutes per unit since otherwise Emily’s process will be the new bottleneck operation with higher processing time and consequently lesser system productivity .

Therefore ,

4 + 60/N < = 10

Or, 60/N < = 6

Or, N > = 10

REQUIRED BATCH SIZE = 10